Use the table below to answer the following questions:
Present Value of 1 Factor Present Value of an Annuity of 1 Factor
Period 1/2 Yr Full-Yr 1/2 Yr Full-Yr
1 0.9449 0.8929 0.9449 0.8929
2 0.8929 0.7972 1.8378 1.6901
3 0.8437 0.7118 2.6814 2.4018
4 0.7972 0.6355 3.4786 3.0373
5 0.7533 0.5674 4.2319 3.6048
6 0.7118 0.5066 4.9437 4.1114
Assumption: Required annual effective rate (EPR) of return is 12%.

If an investment pays you $8,500 every 6 months for 3 years, starting at the beginning of each 6 month period, what is its present value?

Group of answer choices

44,471

44,021

40,831

36,302
Answer #1

Number of payments = 3*2 = 6

Present value = amount to be received today + present value of amount to be received in future

= 8500 +8500*PVAF(1/2 Year, 5 periods)

=8500+8500*4.2319

=$44,471.15

I.e. $44,471